OpenStudy (anonymous):

Given that sin 18 (deg) = (sqrt(5)-1)/4 ... find exact expression for cos 36 (deg). **cos(2x) = 1-2sin^2x

OpenStudy (freckles):

well 2*18=36

OpenStudy (freckles):

plug into the identity

OpenStudy (freckles):

replace x in cos(2x)=1-2sin^2(x) with 18

OpenStudy (anonymous):

I don't get the step where 1-2((sqrt(5)-1)/4)^2 = 1-2((3-sqrt(5))/8)

OpenStudy (freckles):

\[\cos(2 \cdot 18)=\cos(36)=1-2\sin^2(18) =1-2(\sin(18))^2 \\ =1-2 (\frac{\sqrt{5}-1}{4})^2\]

OpenStudy (freckles):

\[=1-2 \frac{(\sqrt{5}-1)^2}{4^2}=1-\frac{2}{16}(\sqrt{5}-1)^2=1-\frac{1}{8}(\sqrt{5}-1)^2\]

OpenStudy (freckles):

multiply the following: \[(\sqrt{5}-1)^2=(\sqrt{5}-1)(\sqrt{5}-1)=?\]

OpenStudy (anonymous):

solution : (sqrt(5) + 1) / 4

OpenStudy (freckles):

that is right

OpenStudy (freckles):

did you get that or is that the answer in the back?

OpenStudy (anonymous):

that's what it says in the back..

OpenStudy (freckles):

ok do you still want help on this...

OpenStudy (freckles):

I was trying to get you to multiply (Sqrt(5)-1)(sqrt(5)-1)

OpenStudy (anonymous):

yes

OpenStudy (freckles):

\[(\sqrt{5}-1)^2 \\ =(\sqrt{5}-1)(\sqrt{5}-1) \\ =\sqrt{5}(\sqrt{5}-1)-1(\sqrt{5}-1)=?\]

OpenStudy (anonymous):

\[6 - \sqrt{5}-\sqrt{5}\]

OpenStudy (freckles):

\[(\sqrt{5}-1)^2=5-\sqrt{5}-\sqrt{5}+1=6-2 \sqrt{5}\] good so we have \[\cos(2 \cdot 18)=\cos(36)=1-2\sin^2(18) =1-2(\sin(18))^2 \\ =1-2 (\frac{\sqrt{5}-1}{4})^2 =1-2(\frac{6- 2 \sqrt{5}}{16})=1-\frac{6-2 \sqrt{5}}{8} \\ =\frac{8}{8}-\frac{6-2 \sqrt{5}}{8}=\frac{8-(6-2 \sqrt{5})}{8}\]

OpenStudy (freckles):

simplify the top

OpenStudy (freckles):

\[8-(6-2 \sqrt{5})=8-6+2\sqrt{5}=?\]

OpenStudy (anonymous):

4\[4\sqrt{5}\]

OpenStudy (freckles):

well 8-6 is 2 so \[8-(6-2 \sqrt{5})=8-6+2\sqrt{5}=2+2 \sqrt{5}\]

OpenStudy (freckles):

you can't do anything else with the top except factor a 2 out like so \[2(1+\sqrt{5})\]

OpenStudy (freckles):

so we have \[\cos(2 \cdot 18)=\cos(36)=1-2\sin^2(18) =1-2(\sin(18))^2 \\ =1-2 (\frac{\sqrt{5}-1}{4})^2 =1-2(\frac{6- 2 \sqrt{5}}{16})=1-\frac{6-2 \sqrt{5}}{8} \\ =\frac{8}{8}-\frac{6-2 \sqrt{5}}{8}=\frac{8-(6-2 \sqrt{5})}{8} \\ =\frac{2(1+\sqrt{5})}{8}\]

OpenStudy (anonymous):

and then divide top and bottom by 2

OpenStudy (freckles):

and we know 2/8 reduces to?

OpenStudy (freckles):

yea

OpenStudy (freckles):

so you would have what you seek

OpenStudy (anonymous):

would ya look at that.

OpenStudy (anonymous):

who woulda thunk it

OpenStudy (freckles):

a little bit of algebra but does everything I asked you to do and everything I did make sense?

OpenStudy (freckles):

do you have any question about what I did?

OpenStudy (anonymous):

no i got it <3

OpenStudy (anonymous):

thanks

OpenStudy (freckles):

cool